# FauForensics: Boosting Audio-Visual Deepfake Detection with Facial Action Units

### v1

@@ Abstract
The rapid evolution of generative AI has increased the threat of realistic audio-visual deepfakes, demanding robust detection methods. Existing solutions primarily address unimodal (audio or visual) forgeries but struggle with multimodal manipulations due to inadequate handling of heterogeneous modality features and poor generalization across datasets. To this end, we propose a novel framework called FauForensics by introducing biologically invariant facial action units (FAUs), which is a quantitative descriptor of facial muscle activity linked to emotion physiology. It serves as forgery-resistant representations that reduce domain dependency while capturing subtle dynamics often disrupted in synthetic content. Besides, instead of comparing entire video clips as in prior works, our method computes fine-grained frame-wise audiovisual similarities via a dedicated fusion module augmented with learnable cross-modal queries. It dynamically aligns temporal-spatial lip-audio relationships while mitigating multi-modal feature heterogeneity issues. Experiments on FakeAVCeleb and LAV-DF show state-of-the-art (SOTA) performance and superior cross-dataset generalizability with up to an average of 4.83% than existing methods.
@@ I Introduction
Compared with unimodal deepfakes, advanced multimodal forgeries can bring a more realistic experience, which raises increasing security concerns about multimodal deepfakes in the community. Recently, researchers have proposed some detectors to capture multimodal forgery traces and obtain promising performance. However, increasingly advanced and various multimodal forgery attacks bring new challenges and demands to current detectors, especially accurately and simultaneously capturing various forgeries including audio forgery, visual forgery, and multi-modal forgery. Therefore, it is urgent to develop a powerful and robust multi-task deepfake detector against various audio-visual forgery attacks.
@@
Previous audio-visual deepfake detectors usually focus on inter-modal dissonance [2], ensemble learning [3], and multimodal joint learning [4] to capture audio-visual forgery traces. For example, Chugh et al. [2] computed modality dissonance score between audio and visual features via a contrastive learning method for detecting multimodal forgery, which lacks inter-modal feature interaction. Hashmi et al. [3] utilize an ensemble-based method to distinguish the authenticity of the audio-video pair. In summary, these methods either lack inter-modal or intra-modal interaction, which limits the detector’s ability to capture various forgeries including audio forgery, visual forgery, and multi-modal forgery. Subsequently, some approaches [5, 6, 7] attempt to focus on both inter- and intra-modal interaction, but these methods are only used for detecting multi-modal forgery and not for unimodal forgery. Recently, a few work [8, 9] have paid attention to this problem and proposed to capture the above three forgeries simultaneously. Although these methods have gained promising performance under within-database settings, most of them easily suffer from over-fitting when handling unseen audio-visual signals.
@@
To address these challenges, we present FauForensics, an audio-visual deepfake detection framework guided by Facial Action Units (FAUs) to identify subtle manipulation patterns and enhance generalization. Our key insight stems from analyzing the temporal correlation intensity of FAUs: real videos exhibit significantly higher consistency in FAU-based muscle movements compared to forged ones (Fig. 1(b)), where discontinuous facial dynamics reveal cross-modal dissonance. The framework operates through four stages (see Fig. 2): FAUs-enhanced feature learning: A frozen pre-trained FAU encoder extracts fine-grained micro-motion features, while a trainable video encoder captures global visual context. Implicit feature alignment: A query-shared multimodal transformer implicitly aligns audio-visual representations via attention mechanisms. Temporal attentional pooler: We introduce a novel frame-wise temporal attentional pooling module to jointly capture intra-modality inconsistencies (e.g., irregular eye blinks) and inter-modality conflicts (e.g., mismatched lip-audio sync). Multiple forgery detection: Pooled features are processed through modality-specific MLPs to predict authenticity scores for unimodal/multimodal scenarios. This design shifts focus from clip-level coarse comparisons to frame-level spatiotemporal anomaly detection driven by physiologically grounded FAUs, significantly improving robustness against evolving forgery techniques
@@
  • Experiments on FakeAVCeleb and LAVDF show state-of-the-art (SOTA) performance. Importantly, it demonstrates superior cross-dataset generalizability with up to an average of 4.83% to existing methods.
@@ II-A Audio-Visual Deepfake Detection
Most of the prior work focused on unimodal deepfake detection [10, 11, 12], especially image-level forgery detection. With the rapid development of multimodal generative models, some researchers [13, 3, 14, 15, 16, 17] were turning to detect multimodal forgery traces using audio-visual signals. For example, Agarwal et al. [18] claimed that the inconsistency between the dynamics of the mouth shape (i.e., visemes) and the spoken phoneme could be used to capture manipulation traces. Although they used hand-crafted features to detect audio-visual forgery traces, they ignored the feature interaction between audio-visual signals. To further mine complementary information between modalities, Zhou et al. [13] proposed a joint detection framework to learn robust feature representations from audio-visual signals. These methods motivate us to detect deepfakes using multi-modal signals. Therefore, in this work, we focus on multimodal deepfake detection and propose a novel framework to model inter- and intra-modal relationships between audio-visual signals.
@@ II-B Facial Attribution-Guided Forgery Detection
To better capture audio-visual forgery traces, some researchers attempted to capture multimodal forgery traces from basic facial characteristics. For example, Mittal et al. [19] mined emotion cues from the two modalities and detected artifacts by emotion similarity (or dissimilarity) between audio-visual signals. Although these prior works obtained promising performance in within-database evaluations, they easily suffer from over-fitting and do not cope with the increasingly advanced and various multimodal forgery attacks. To avoid over-fitting problems, recent work [20, 4, 21, 7] attempted to employ self-supervised learning methods pre-trained in a large-scale authentic database to mine facial-related inherent properties, such as identity consistency, lip sync, etc. However, these large-scale pre-training methods are time-consuming and resource-hungry. Recently, some visual forgery detectors [22, 23] utilized facial action units (FAUs) to capture visual forgery traces related to facial muscle but neglected to exploit the correlation between audiovisual modalities. Facial Action Units (FAUs) exhibit strong correlations with audio signals through both physiological relationships (e.g., lower-face AUs drive different sound generation) and affective synchronization (e.g., eyebrow movements are correlated with voice frequency) [24]. Therefore, we introduce FAUs to obtain structured geometric information for better capturing audio-visual dissonance between facial muscle movements and audio signals.
@@ II-C Temporal Forgery Detection
Previously, some works [25, 26, 27] attempted to capture temporal forgery cues for unimodal deepfake detection. However, these methods relay intra-modal modeling of unimodality in the temporal domain and ignore inter-modal temporal forgery traces for multimodal deepfake detection. To capture inter-modality temporal forgery traces, most research efforts were devoted to exploring audio-visual dissonance [28, 29, 30, 31] and joint audio-visual feature learning [13, 5, 7, 32] for improving generalization across unseen manipulation methods. For example, Zhou et al. [13] proposed a multimodal joint learning approach to learn robust feature representations from audio-visual modality. Liu et al. [5] proposed a contrastive learning method to reduce the cross-modal gap and learn a compositional embedding from multimodal signals for capturing audio-visual temporal forgeries. However, these methods focused on clip-level or patch-level semantic interaction between two modalities and ignored key temporal correlations for audio-visual temporal forgery detection. To avoid these issues, a few researchers [9] began to model more fine-grained inter-modality relationships via frame-level feature interaction to capture audio-visual dissonance. For example, Yin et al. [9] proposed a graph attention network to learn audio-visual synchronization patterns for capturing audio-visual forgery traces. However, this method relied on randomly sampled frames from video clips to learn clip-level feature interaction between two modalities, which ignored more fine-grained frame-by-frame forgery clues. In this work, we explore intra- and inter-modality temporal correlation in a frame-by-frame manner to capture slight temporal forged artifacts for audio-visual temporal forgery detection.
Figure 2: Overview of the proposed FauForensics framework. Firstly, we take audio-video pairs as inputs and encode inputs into latent features via two audio/video encoders with learnable weights and a pre-trained FAU encoder with frozen weights. FAU features are expected to enable detectors to capture FAU-related forgeries. We then perform implicit feature alignment before inter-modal modeling for better feature interaction. Next, we design a novel temporal attentional pooler to model intra- and inter-modal temporal correlation for capturing various forgeries. Finally, we feed pooled features into MLP for classification.
@@ III-A Architecture
Our proposed FauForensics is a novel audio-visual deepfake detection architecture that aims to learn robust and fine-grained temporal consistency between audio-visual signals. Our framework consists of FAU-enhanced Feature Learning, Implicit Feature Alignment, Temporal Attentional Pooler, and Multiple Forgery Detection. We depict an overview of our proposed approach in Fig. 2. We can clearly see that three different encoders (i.e., $f^{\rm a}_{\rm enc}$, $f^{\rm v}_{\rm enc}$, $f^{\rm au}_{\rm enc}$) take audio and visual signals as inputs and output corresponding latent representations, respectively. In particular, we introduce a facial action units (FAU) encoder to capture fine-grained facial motion based facial structured geometric information from video signals. To better learn information interaction between modalities, we first perform feature alignment across the temporal domain. Then, we propose a temporal attentional pooler in a frame-by-frame manner to capture temporal consistency via inter- and intra-modal modeling. Finally, we feed flattened pooled vectors into MLP-based classifiers for unimodal/multimodal forgery predictions.
@@ III-B FAU-enhanced Feature Learning
For the visual modality, we first extract $N$ consecutive video frames ${\bm{X}}_{\rm v}\in\mathbb{R}^{T\times C\times H\times W}$ from the given video. To better enable the model to focus on fine-grained facial motion across the temporal domain, we utilize the pre-trained FAU encoder $f^{\rm au}_{\rm enc}$ with frozen weights for FAU feature extraction. FAU features can enable forgery detectors to capture FAU-related forgeries, which can enhance visual representations. The FAU latent features are defined as ${\bm{Z}}_{\rm au}=f_{\rm enc}^{\rm au}({\bm{X}}_{\rm v})$. Moreover, to learn richer visual representations, we use the video encoder $f^{\rm v}_{\rm enc}$ with learnable weights for further extracting video features ${\bm{Z}}_{\rm vid}=f_{\rm enc}^{\rm v}({\bm{X}}_{\rm v})$. Finally, we fuse the two features via a simple addition operation $f_{\rm add}$ and obtain FAU-enhanced visual latent features ${\bm{Z}}_{\rm v}=f_{\rm add}({\bm{Z}}_{\rm vid},{\bm{Z}}_{\rm au})$, which has the same size as the audio latent features.
To better analyze audio signals, we extract mel-spectrogram from raw audio waveforms. The resulting Mel-spectrogram $\bm{X}_{\rm a}$ is a 2-dimensional representation of sound signals. Therefore, Mel-spectrogram is essentially a single-channel image including time and frequency information simultaneously, which can be taken as the input of audio encoder $f^{\rm a}_{\rm enc}$. The resulting audio latent features ${\bm{Z}}_{\rm a}\in\mathbb{R}^{T\times L}$ can be denoted as ${\bm{Z}}_{\rm a}=f_{\rm enc}^{\rm a}({\bm{X}}_{\rm a})$.

### v2

@@ Abstract
The rapid evolution of generative AI has intensified the threat of realistic audio-visual deepfakes, demanding robust and generalizable detection methods. Existing solutions primarily address unimodal (e.g., audio, visual) forgeries but struggle with multimodal manipulations due to inadequate handling of heterogeneous modality features and poor cross-dataset generalization. We propose FauForensics, a novel framework leveraging biologically invariant facial action units (FAUs), which are quantitative descriptors of facial muscle activity linked to emotion physiology. They serve as forgery-resistant representations that reduce domain dependency while capturing subtle synthetic-content disruptions. In addition, unlike prior clip-level comparisons, our method computes frame-wise audio-visual similarities via a fusion module with learnable cross-modal queries, dynamically aligning lip-audio relationships and mitigating feature heterogeneity. Experiments on FakeAVCeleb and LAV-DF show state-of-the-art performance with 4.83% average cross-dataset improvement over existing methods.
Index Terms:
Deepfake Detection, Multimodal, Facial Action Units, Implicit Feature Alignment, Temporal Correlation.
@@ I Introduction
Compared with unimodal deepfakes, advanced multimodal forgeries can bring a more realistic experience, which raises increasing security concerns about multimodal deepfakes in the community. Recently, researchers have proposed some audio-visual deepfake detectors to capture multimodal forgery traces and obtain promising performance. However, increasingly advanced and various multimodal forgery attacks bring new challenges and demands to current detectors, especially accurately and simultaneously capturing various forgeries, including audio forgery, visual forgery, and multimodal forgery. Therefore, it is urgent to develop a powerful and robust multi-task deepfake detector against various audio-visual forgery attacks.
@@
Previous audio-visual deepfake detectors usually focus on inter-modal dissonance [2], ensemble learning [3, 4], and multimodal joint learning [5] to capture audio-visual forgery traces. For example, Chugh et al. [2] computed modality dissonance score between audio and visual features via a contrastive learning method for detecting multimodal forgery, which lacks inter-modal feature interaction. Hashmi et al. [3] utilize an ensemble-based method to distinguish the authenticity of the audio-video pair. In summary, these methods either lack inter-modal or intra-modal interaction, which limits the detector’s ability to comprehensively capture various forgeries, including (1) audio-only forgeries (e.g., synthetic voice generation, audio splicing), (2) visual-only forgeries (e.g., face swap), and (3) multi-modal forgeries (e.g., misaligned lip-sync, mismatched audio-visual events). Subsequently, some approaches [6, 7, 8] attempt to focus on both inter- and intra-modal interaction, but these methods are only used for detecting multi-modal forgeries and not for unimodal forgeries. Recently, a few work [9, 10] have paid attention to this problem and proposed methods to simultaneously capture all three types of forgeries. Although these methods have gained promising performance under within-database settings, most of them easily suffer from obvious performance degradation when handling unseen audio-visual signals.
To address these challenges, it is essential to develop a generalized and robust detection framework capable of effectively capturing intrinsic unimodal/multimodal forgery traces for accurate detection. One promising solution is to transfer intrinsic knowledge learned from real audio-visual signals via large-scale multimodal pretraining to downstream detection tasks. For example, Oorloff et al. [8] propose a two-stage cross-modal learning method to capture the intrinsic audio-visual correspondences via self-supervision on large-scale real videos. However, these pre-training methods rely on multi-stage training with large-scale datasets are time-consuming and resource-hungry. Furthermore, cross-modal biological invariance between modalities presents another possible solution for multimodal forgery detection. For example, early work [11] utilizes emotion consistency to capture affective forgery cues between audio-visual signals. However, emotion-only forgery detectors can not handle increasingly improved multimodal deepfakes. Therefore, it is essential to capture more effective physiologically constrained relationships between modalities, as these are inherently difficult to perfectly replicate in synthetic multimedia content.
@@
In this work, we present FauForensics, an audio-visual deepfake detection framework guided by Facial Action Units (FAUs) to identify subtle manipulation patterns and enhance generalization. Our key insight stems from analyzing the temporal correlation intensity of FAUs: real videos exhibit significantly higher consistency in FAU-based muscle movements compared to forged ones (Fig. 1(b)), where discontinuous facial dynamics reveal cross-modal dissonance. The framework operates through four stages (see Fig. 2): FAU-enhanced feature learning: A frozen pre-trained FAU encoder extracts fine-grained micro-motion features, while a trainable video encoder captures global visual context. Implicit feature alignment: A query-shared multimodal transformer implicitly aligns audio-visual representations via attention mechanisms. Temporal attentional pooler: We introduce a novel frame-wise temporal attentional pooling module to jointly capture intra-modality inconsistencies (e.g., irregular eye blinks) and inter-modality conflicts (e.g., mismatched lip-audio sync). Multiple forgery detection: Pooled features are processed through modality-specific MLPs to predict authenticity scores for unimodal/multimodal scenarios. This design shifts focus from clip-level coarse comparisons to frame-level spatiotemporal anomaly detection driven by physiologically grounded FAUs, significantly improving robustness against evolving forgery techniques.
@@
  • Experiments on FakeAVCeleb and LAV-DF show state-of-the-art (SOTA) performance. Importantly, it demonstrates superior cross-dataset generalizability with up to an average of 4.83% to existing methods.
The rest of the paper is organized as follows: Sec. II reviews related work of Audio-Visual Deepfake Detection, Facial Attribution-Guided Forgery Detection, and Temporal Forgery Detection. The proposed FauForensics architecture is detailed in Sec. III. Experimental setup, results, and analysis are described in Sec. IV. Finally, the paper is concluded in Sec. V.
@@ II Related Work
In this section, we give a brief review of Audio-Visual Deepfake Detection, Facial Attribution-Guided Forgery Detection, and Temporal Forgery Detection.
@@ II-A Audio-Visual Deepfake Detection
Most of the prior work focused on unimodal deepfake detection [12, 13, 14, 15, 16, 17], especially visual-only deepfake detection. For example, Yang et al. [18] proposed a lip-based speaker authentication system that focuses on dynamic lip movement to defend against human imposters. With the rapid development of multimodal generative models [19, 20, 21], some researchers [22, 3, 23, 24, 25, 26, 27] were turning to detect multimodal forgery traces using audio-visual signals. For example, Agarwal et al. [28] claimed that the inconsistency between the dynamics of the mouth shape and the spoken phoneme could be used to capture cross-modal manipulation traces. Although they achieved promising performance using hand-crafted features to capture audio-visual forgery traces, they overlook high-level semantic information and the interaction between audio-visual features. To further mine complementary information between modalities, Zhou et al. [22] proposed a joint detection framework to learn robust feature representations from audio-visual signals. In addition, Zou et al. [7] employed modality-specific and cross-modality regularization to improve multimodal representation learning for audio-visual deepfake detection. These methods inspire us to detect subtle cross-modal manipulation traces using multi-modal signals for audio-visual deepfake detection. Compared with unimodal deepfake detection, multimodal deepfake detection approaches require more inter-modal correlation learning to capture complex dependencies across different modalities. Therefore, in this work, we focus on audio-visual deepfake detection and propose a novel framework to model inter- and intra-modal relationships between audio and visual signals.
@@ II-B Facial Attribution-Guided Forgery Detection
To better capture audio-visual forgery traces, some researchers attempted to capture multimodal forgery traces from basic facial characteristics. For example, Mittal et al. [11] mined emotion cues from the two modalities and detected artifacts by emotion similarity (or dissimilarity) between audio-visual signals. Although these prior works obtained promising performance in within-database evaluations, they easily suffer from performance drop and do not cope with the increasingly advanced and various multimodal forgery attacks. To avoid these problems, recent work [29, 5, 30, 8] attempted to employ self-supervised learning methods pre-trained in a large-scale authentic database to mine facial-related inherent properties, such as identity consistency, lip sync, etc. However, these large-scale pre-training methods are time-consuming and resource-hungry. Recently, Liu et al. [31] exploited the biologically intrinsic correlation between lip movements and head postures to detect lip-syncing deepfakes. Peng et al. [32] proposed a gaze analysis model to capture distribution differences of gaze direction pattern between the real and forged videos. In addition, some visual forgery detectors [33, 34] utilized facial action units (FAUs) to capture visual forgery traces related to facial muscle but neglected to exploit the correlation between audio-visual modalities. Facial Action Units (FAUs) exhibit strong correlations with audio signals through both physiological relationships (e.g., lower-face AUs drive different sound generation) and affective synchronization (e.g., eyebrow movements are correlated with voice frequency) [35]. Therefore, we introduce FAUs to obtain structured geometric information for better capturing audio-visual dissonance between facial muscle movements and audio signals.
Figure 2: Overview of the proposed FauForensics framework. Firstly, we take audio-video pairs as inputs and encode them into latent features using two audio/video encoders with learnable weights, alongside a pre-trained FAU encoder with frozen weights. The FAU features are expected to enable detectors to capture FAU-related forgeries. We then perform implicit feature alignment before inter-modal modeling to facilitate more effective audio-visual feature interaction. Next, we design a novel temporal attentional pooler to model intra- and inter-modal temporal correlation for capturing various forgeries. Finally, we feed pooled features into MLP for classification.
@@ II-C Temporal Forgery Detection
Previously, some work [36, 37, 38, 39] attempted to capture temporal forgery cues for unimodal deepfake detection. However, these methods rely on intra-modal modeling of unimodality in the temporal domain and ignore inter-modal temporal forgery traces for multimodal deepfake detection. To capture inter-modality temporal forgery traces, most research efforts were devoted to exploring audio-visual dissonance [40, 41, 42, 31, 43] and joint audio-visual feature learning [44, 8, 45, 46] for improving generalization across unseen manipulation methods. For example, Zhou et al. [22] proposed a multimodal joint learning approach to learn robust feature representations from audio-visual modality. Liu et al. [6] proposed a contrastive learning method to reduce the cross-modal gap and learn a compositional embedding from multimodal signals for capturing audio-visual temporal forgeries. However, these methods focused on clip-level or patch-level semantic interaction between two modalities and ignored key temporal correlations for audio-visual temporal forgery detection. To avoid these issues, a few researchers [10] began to model finer-grained inter-modality relationships via frame-level feature interaction to capture audio-visual dissonance. For example, Yin et al. [10] proposed a graph attention network to learn audio-visual synchronization patterns for capturing audio-visual forgery traces. However, this method relied on randomly sampled frames from video clips to learn clip-level feature interaction between two modalities, which ignored finer-grained frame-by-frame forgery clues. In this work, we explore intra- and inter-modality temporal correlation in a frame-by-frame manner to capture slight temporal artifacts for audio-visual temporal forgery detection.
@@ III Method
In this section, we first introduce FauForensics architecture in Sec.III-A and then present FAU-enhanced Feature Learning, Implicit Feature Alignment, Temporal Attentional Pooler, and Multiple Forgery Detection in Sec. III-B, Sec. III-C, Sec.III-D, and Sec.III-E, respectively. Finally, we describe our implementation details in Sec.III-F.
@@ III-A Architecture
Our proposed FauForensics is a novel audio-visual deepfake detection architecture that aims to learn robust and fine-grained temporal consistency between audio-visual signals. Our architecture consists of FAU-enhanced Feature Learning, Implicit Feature Alignment, Temporal Attentional Pooler, and Multiple Forgery Detection. We depict an overview of our proposed approach in Fig. 2. We can see that three different encoders (i.e., $f^{\rm a}_{\rm enc}$, $f^{\rm v}_{\rm enc}$, $f^{\rm au}_{\rm enc}$) take audio and visual signals (i.e., ${\bm{X}}_{\rm a}$ and ${\bm{X}}_{\rm v}$) as inputs and output corresponding latent representations, respectively. In particular, we introduce a facial action unit (FAU) encoder $f^{\rm au}_{\rm enc}$ designed to extract fine-grained structured geometric information from facial motions in video streams. Then, we apply a fusion operation $f_{\rm\varphi}$ to derive FAU-enhanced visual representations by integrating video features with FAU features. To better learn information interaction between modalities, we first perform implicit feature alignment across the temporal domain via our designed Query-shared audio/visual transformers $f_{\rm at}^{\rm q}$/$f_{\rm vt}^{\rm q}$. Then, we propose a temporal attentional pooler $f_{\rm\phi}$ in a frame-by-frame manner to capture temporal consistency via inter- and intra-modal modeling. Finally, we feed flattened pooled vectors into MLP-based classifiers $f_{\rm mlp}$ for audio-visual forgery predictions $\bm{s}$. The process described above can be formally expressed as follows:
${\bm{s}}=f_{\rm mlp}(f_{\rm\phi}(f_{\rm at}^{\rm q}(f_{\rm enc}^{\rm a}({\bm{X}}_{\rm a})),f_{\rm vt}^{\rm q}(f_{\rm\varphi}(f_{\rm enc}^{\rm v}({\bm{X}}_{\rm v}),f_{\rm enc}^{\rm au}({\bm{X}}_{\rm v}))))).$ (1)
@@ III-B FAU-enhanced Feature Learning
For the visual modality, we first extract $T$ consecutive video frames ${\bm{X}}_{\rm v}\in\mathbb{R}^{T\times C\times H\times W}$ from the given video. To better enable the model to focus on fine-grained facial motion across the temporal domain, we utilize the pre-trained FAU encoder $f^{\rm au}_{\rm enc}$ with frozen weights for FAU feature extraction. FAU features can help forgery detectors capture FAU-related forgeries, thereby enhancing visual representation learning for audio-visual deepfake detection. The FAU latent features are obtained as ${\bm{Z}}_{\rm au}=f_{\rm enc}^{\rm au}({\bm{X}}_{\rm v})$. Moreover, to learn richer visual representations, we use the video encoder $f^{\rm v}_{\rm enc}$ with learnable weights for further extracting video features ${\bm{Z}}_{\rm vid}=f_{\rm enc}^{\rm v}({\bm{X}}_{\rm v})$. Finally, we utilize a feature fusion operation to obtain FAU-enhanced visual latent features ${\bm{Z}}_{\rm v}=f_{\rm\varphi}({\bm{Z}}_{\rm vid},{\bm{Z}}_{\rm au})$, which are dimensionally aligned with the audio latent features ${\bm{Z}}_{\rm a}\in\mathbb{R}^{T\times L}$.
To better analyze audio signals, we extract a mel-spectrogram from raw audio waveforms. The resulting Mel-spectrogram $\bm{X}_{\rm a}$ is a 2-dimensional representation of sound signals. Therefore, Mel-spectrogram is essentially a single-channel image including time and frequency information simultaneously, which can be taken as the input of the audio encoder $f^{\rm a}_{\rm enc}$. The resulting audio latent features ${\bm{Z}}_{\rm a}$ can be denoted as ${\bm{Z}}_{\rm a}=f_{\rm enc}^{\rm a}({\bm{X}}_{\rm a})$.